# On the Jacobs-de Leeuw-Glicksberg decomposition

### v1

@@ 4 The almost weakly stable part of the JdLG-decomposition
We begin with a characterization of the almost weakly stable part of the JdLG-decomposition through the unique invariant mean on the space of weakly almost periodic functions. This characterization is related to the so-called ”flight functions" discussed in [18, 30, 40].
@@
    There exist $p>0$ such that for all $f^{\prime}\in E^{\prime}$, we have
    $m\left(s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\right)=0.$ (6)
@@
We will first show that (i) implies (ii). Let $(s_{i})_{i\in I}\subseteq S$ be a sequence for which $\lim_{i}\pi_{s_{i}}\xi=0$ weakly, and let $p>0$ be arbitrary. Since for each $f^{\prime}\in E^{\prime}$ we have $s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\in\text{W}(G)$ we have that $\{|f^{\prime}(\pi_{ss_{i}}\xi)|^{p}\}_{n=1}^{\infty}$ is a relatively weakly compact set, so there exists a subnet $(s_{j})_{j\in J}$ for which $|f^{\prime}(\pi_{ss_{j}}\xi)|^{p}$ converges weakly to some function $\phi(s)$. Since weak convergence implies pointwise convergence and $\lim_{j}\pi_{s_{j}}\xi=0$ weakly, we see that $\phi$ is the 0 function, hence
@@
It is clear that (ii) implies (iii), so we will now show that (iii) implies (i). Let us assume that $\xi\notin E_{\text{aws}}$ and let us first consider the case of $p=2$. Since $\pi$ is a JdLG-admissible representation of $S$, can write $\xi=\xi_{1}+\xi_{2}$ with $0\neq\xi_{1}\in E_{\text{rev}}$ and $\xi_{2}\in E_{\text{aws}}$. Since $E_{\text{rev}}$ can be further decomposed into a direct sum of irreducible finite dimensional subspaces, let us write $\xi_{1}=x_{1}+x_{2}$ with $x_{1}$ belonging to a finite dimensional irreducible subspace $E_{1}$. We may further assume without loss of generality that $S$ acts on $E_{1}$ by unitary matrices. If $\text{dim}(E_{1})=k$, then for each $v\in E_{1}$ there exists $k$ functionals $f_{1},\cdots,f_{k}$ (depending on $v$) for which $\|v\|^{2}=\sum_{j=1}^{k}|f_{j}(\pi_{s}v)|^{2}$ for all $s\in S$. It follows that
$\|x_{1}\|^{2}=m\left(\|\pi_{s}x_{1}\|^{2}\right)=m\left(\sum_{j=1}^{k}|f_{j}(\pi_{s}x_{1})|^{2}\right)=\sum_{j=1}^{k}m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right),$ (8)
so for some $1\leq j\leq k$ we have
$m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right)\geq\frac{\|x_{1}\|^{2}}{k}>0.$ (9)
We extend $f_{j}$ to a functional $\varphi$ on all of $E$ by defining $\varphi(v)=0$ for all $v$ in the complement of $E_{1}$, and we see that
$m\left(|\varphi(\pi_{s}\xi)|^{2}\right)>0.$ (10)
To handle the case of a general $p>0$, we let $S^{w}$ denote the weakly almost periodic compactification of $S$, and we let $i,M,$ and $\mu$ be as in Section 2.1. Let $f^{\prime}\in E^{\prime}$ be arbitrary. Since
$m\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)=M\left(i\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)\right)=\int_{S^{w}}\left|f^{\prime}(\pi_{s}\xi)\right|^{p}d\mu,$ (11)
we see that $m(|f^{\prime}(\pi_{s}\xi|^{p})=0$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e. $s\in S^{w}$. Lastly, we observe that $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e. $s\in S^{w}$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{2}=0$ for $\mu$-a.e. $s\in S^{w}$. ∎
Our next result can be seen as a generalization of [21, Theorem 16.34] (see also [31] and [22, Theorem 2.5]), and also relates the almost weakly stable vectors to the vectors from $\mathcal{H}_{w}$ in the compact-weak mixing decomposition.
@@
    $d_{*}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
@@
    $d_{\mathcal{F}}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
@@
    $\underline{d}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
@@
If we assume even further that $E^{\prime}$ is separable and $\lim_{n\rightarrow\infty}\lambda(F_{n})=\infty$, then (i)-(ix) are equivalent.
@@
    There exists a set $A\subseteq S$ with $d_{\mathcal{F}}(A)=1$ such that $\lim_{\underset{s\in A}{s\to\infty}}T_{s}\xi=0$ weakly.
@@
$E_{\text{aws}}$ $=\left\{\xi\in E\ |\ m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}\right\},$
@@
To see that (iii) implies (ii), we observe that
@@
It is clear that (iv) implies (iii). To see that (iii) implies (iv), we recall that (iii) is equivalent to (i), and (i) is independent of the choice of $m$. In particular, if (iii) holds for a particular left-invariant mean, then (iii) holds for all left-invariant means, which yields (iv).
Now let us assume that $(S,\lambda)$ admits a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$. We observe that $g(s):=|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}$ is in $W(S)$. For each $\in\mathcal{M}(\mathcal{F}), mm$ restricts to the unique bi-invariant mean on $W(s)$, so the value of $m(g)$ is independent of the choice of $m$, hence
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=m(g)\text{ for all }m\in\mathcal{M}(\mathcal{F}).$ (14)
In particular, the limit in the left hand side of (14) exists. Recalling that if (ii) holds for a particular left-invariant mean then it holds for all left-invariant means, we see that (ii) is equivalent to (v). Similarly, we see that (iii) is equivalent to (vi).
It is clear that (vii) implies (vi). To see that (vi) implies (vii), it suffices to observe that (vi) is equivalent to (i), so if (vi) holds for a particular left-Følner sequence $\mathcal{F}$, then (vi) holds for every left-Følner sequence $\mathcal{F}$.
Now let us assume that $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence. It is now clear that (viii) implies (v), so we will proceed to show that (v) implies (viii). Since $(\pi_{s})_{s\in S}$ is uniformly bounded, we suppose without loss of generality that $S$ is a monoid and that $(\pi_{s})_{s\in S}$ consists of contractions. Then the adjoint action $(\pi_{s}^{*})_{s\in S}$ leaves the ($\sigma^{*}$-compact) dual unit ball $D^{\prime}\subseteq E^{\prime}$ invariant. Let $(T_{s})_{s\in S}$ denote the Koopman action on $C(D^{\prime})$ induced by $(\pi^{*}_{s})_{s\in S}$. For $\xi$ satisfying (ii) we let $f_{\xi}\in C(D^{\prime})$, $f_{\xi}(x^{\prime}):=\left|\langle\xi,x^{\prime}\rangle\right|^{p}$, $p>0$. We observe by assumption, that
@@
holds pointwise. Then, by the dominated convergence theorem the convergence of 15 holds weakly, i.e.,
@@
holds for any $\mu\in\left(C(D^{\prime})\right)^{\prime}$. By Eberlein’s generalized mean ergodic theorem, weak convergence implies the convergence of 16 in norm (see, e.g., Chapter 2, Theorem 1.5 of [34]). Hence
@@
Now let us show that (vi) is equivalent to (ix) under additional stated assumptions. Since $d_{\mathcal{F}}(C)=0$ for every compact set $C$, we see that (ix) implies (vi), so it remains to show that (vi) implies (ix). Since $E^{\prime}$ is separable, it is first countable, so let $\{U_{n}\}_{n=1}^{\infty}$ be a neighborhood basis for $0\in E^{\prime}$. For each $n\in\mathbb{N}$, let $A_{n}=\{s\in S|\pi_{s}\xi\in U_{n}\}$ and let $m_{n}$ be such that $\lambda(A_{n}\cap F_{m})>\frac{n-1}{n}\lambda(F_{m})$ for all $m\geq m_{n}$. Assuming without loss of generality that $m_{n+1}>m_{n}$, we see that it suffices to take $A=\bigcup_{n=1}^{\infty}\bigcup_{m=m_{n}+1}^{m_{n+1}}(A_{n}\cap F_{m})$. ∎
@@
We now recall [32, Example 4.3] to give an example showing that Lemma 4.1 does not necessarily hold if $W(S)$ only admits a left-invariant mean and not a bi-invariant mean. This example will also show that Theorem 4.2 does not necessarily hold if $S$ is a left-amenable subsemigroup of an amenable group that is not right amenable.
We consider the semigroup $(S,\cdot)$ given by $S=\{(m,n)\in\mathbb{Z}^{2}\ |\ m\geq 0,n\geq 1\}$ and $(m_{1},n_{1})\cdot(m_{2},n_{2})=(2^{n_{2}}m_{1}+m_{2},n_{1}+n_{2})$. $S$ is cancellative and left-amenable, hence it satisfies the strong Følner condition and possesses a left-Følner sequence $\mathcal{F}=(F_{i})_{i=1}^{\infty}$. Since $S\cdot(0,1)\cap S\cdot(1,1)=\emptyset$, we see that $S$ is not right amenable. Furthermore, by replacing $F_{n}$ with $F_{n}\cdot(0,1)$, we may assume without loss of generality that $F_{n}\subseteq S\cdot(0,1)$ for all $n$. We observe that $S=S\cdot(0,1)\cup S\cdot(1,1)\cup\mathbb{Z}_{\geq 0}\times\{0\}$. Now consider the function $f:S\rightarrow\{0,1\}$ given by $f(s)=1$ if and only if $s\in S\cdot(0,1)$. We see that $\{R_{s}f\}_{s\in S}=\{f,0,1,\mathbbm{1}_{A}-f\}$, where $A=S\cdot(0,1)\cup S\cdot(1,1)$, hence it is relatively weakly compact and we see that $f\in W(S)$. Now let $E=W(S)$, $\xi=f$, $\pi=R$, and $x^{\prime}$ denote evaluation at $s_{0}$ for some $s_{0}\in S$. We see that $f\in E_{\text{aws}}$ because $R_{(1,1)}f=0$. However, if $m\in\mathcal{M}(\mathcal{F})$, then
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|d\lambda(s)=m(s\mapsto\left(R_{s}f\right)(s_{0}))=m(1)=1.$ (17)
@@ 5 The JdLG decomposition via ultrafilters
Let $E=\mathcal{H}$ be a Hilbert space, $U$ a unitary operator on $\mathcal{H}$, and $\mathcal{H}=\mathcal{H}_{c}\oplus\mathcal{H}_{w}$ the compact-weak mixing decomposition. There are various results in the literature, such as [8, Theorem 2.2], [10, Theorem 2.25], [35, Theorem 4], and [12, Section 6], that (allow one to) construct the orthogonal projection $P:\mathcal{H}\rightarrow\mathcal{H}_{c}$ as an ultrafilter limit of powers of $U$. Theorem 5.2 is motivated by these previously mentioned results and attempts to generalize them to JdLG-admissible representations on a Banach space $E$. Strictly speaking, we cannot say that Theorem 5.2 generalizes all of the aforementioned results, because they also consider various classes of polynomial subsequences that we do not consider. For the next two results, we freely use the notation and facts from Subsection 2.5.
@@
We first prove (i). Let $h:\beta S\rightarrow\mathscr{S}_{r}=K(\mathscr{S})$ denote the map $h(q)=\displaystyle q-\lim_{s}\pi_{s}\xi$ with convergence taking place in the weak topology of $E_{r}$ (instead of the weak topology of $E$). Since $\pi:S\rightarrow\mathscr{S}$ is a homomorphism with dense image in a semitopological semigroup and $\pi=h\circ e$ on $S$, we see that $h$ a continuous surjective homomorphism. In particular, $h$ sends idempotents to idempotents. Since $\mathscr{S}_{r}=K(\mathscr{S})$ is a compact topological group, so it contains a unique idempotent $Q$. Furthermore, we have already seen that $Q$ acts as the identity on $E_{r}$, so the desired result follows.
We now proceed to prove (ii). Theorem 4.2 tells us that for any weakly open neighborhood $U$ of 0, we have
$d^{*}\left(\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\right)=0\text{, hence }\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\notin q.$ (18)
Since $q$ is an ultrafilter, we see that $\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\in U\right\}\in q$. Since $U$ was arbitrary, we see that $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$. ∎
@@
We begin with the proof of (i). We let $\mathscr{S}=c\ell_{\sigma}(\pi(S))$ and we recall that $\mathscr{S}$ is a compact semitopological semigroup. Let $\phi:\beta S\rightarrow\mathscr{S}$ be given by $\phi(q)=\pi_{q}=q-\lim_{s}\pi_{s}$, and observe that as in the proof of Lemma 5.1, $\phi$ is continuous surjective homomorphism. It follows that $K(\mathscr{S})=\phi(K(\beta S))$. Since $\pi$ is a JdLG-admissible representation, we see that $K(\mathscr{S})$ is a compact topological group with a unique idempotent $Q$, which is the projection onto $E_{r}$. Since $K(\mathscr{S})$ is closed, we see that $\phi\left(\overline{K(\beta S)}\right)=K(\mathscr{S})$. Since $\phi$ is a homomorphism, $\phi$ maps idempotents to idempotents, hence $\phi$ maps all idempotents of $\overline{K(\beta S)}$ to $Q$, which completes the proof of (i).

### v2

@@ 4 The almost weakly stable part of the JdLG-decomposition
We begin with a characterization of the almost weakly stable part of the JdLG-decomposition through the unique invariant mean on the space of weakly almost periodic functions. This characterization is related to the flight functions discussed in [38, Appendix 10].
@@
    There exist $p>0$ such that for all $f^{\prime}\in E^{\prime}$ Equation (8) holds.
@@
We will first show that (i) implies (ii). Let $(s_{i})_{i\in I}\subseteq S$ be a net for which $\lim_{i}\pi_{s_{i}}\xi=0$ weakly, and let $p>0$ be arbitrary. Since for each $f^{\prime}\in E^{\prime}$ we have $s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\in\text{W}(S)$ we have that $\{|f^{\prime}(\pi_{ss_{i}}\xi)|^{p}\}_{i\in I}$ is a relatively weakly compact set, so there exists a subnet $(s_{j})_{j\in J}$ for which $|f^{\prime}(\pi_{ss_{j}}\xi)|^{p}$ converges weakly to some function $\phi(s)$. Since weak convergence implies pointwise convergence and $\lim_{j}\pi_{s_{j}}\xi=0$ weakly, we see that $\phi$ is the 0 function, hence
@@
It is clear that (ii) implies (iii), so we will now show that (iii) implies (i).
Let $S^{w}$ denote the weakly almost periodic compactification of $S$, and we let $i,M,$ and $\mu$ be as in Section 2.1. Let $f^{\prime}\in E^{\prime}$ be arbitrary. Since
$m\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)=M\left(i\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)\right)=\int_{K(S^{w})}\left|f^{\prime}(\pi_{s}\xi)\right|^{p}d\mu,$ (10)
we see that $m(|f^{\prime}(\pi_{s}\xi|^{p})=0$ if and only if $f^{\prime}(\pi_{s}\xi)=0$ for $\mu$-a.e. $s\in K(S^{w})$, and in fact for every $s\in K(S^{w})$ since this is a continuous function. Since $f^{\prime}$ is arbitrary, this implies $\pi_{s}\xi=0$ for $s\in K(S^{w})$.
It is easy to see that $\pi$ extends to a representation of $S^{w}$ with the image in the weak closure of $\pi(S)$; by definition given in (1), we have $\xi\in E_{\text{aws}}$. ∎
Our next result can be seen as a generalization of [29], [20, Theorem 16.34], and [21, Theorem 2.5]. Furthermore, it relates the almost weakly stable vectors to the vectors from $\mathcal{H}_{w}$ in the compact-weak mixing decomposition.
@@
    $d_{*}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
@@
    $d_{\mathcal{F}}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
@@
    $\underline{d}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
@@
If we assume even further that $E^{\prime}$ is (norm) separable and $\lim_{n\rightarrow\infty}\lambda(F_{n})=\infty$, then (i)-(ix) are equivalent.
@@
    There exists a set $A\subseteq S$ with $d_{\mathcal{F}}(A)=1$ such that $\lim_{\underset{s\in A}{s\to\infty}}\pi_{s}\xi=0$ weakly.
@@
$E_{\text{aws}}$ $=\left\{\xi\in E\ :\ m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}\right\},$
@@
To see that (iii) implies (ii), we observe that for every $x^{\prime}\in E^{\prime}$ and $\epsilon>0$
@@
It is clear that (iv) implies (iii). To see that (iii) implies (iv), it suffices to recall that $d_{*}\leq m$ for any left-invariant mean.
Now let us assume that $(S,\lambda)$ admits a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$. We observe that $\phi(s):=|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}$ is in $W(S)$. For each $\in\mathcal{M}(\mathcal{F}), mm$ restricts to the unique bi-invariant mean on $W(S)$, so the value of $m(\phi)$ is independent of the choice of $m$, hence
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=m(\phi)\text{ for all }m\in\mathcal{M}(\mathcal{F}).$ (13)
In particular, the limit in the left hand side of (13) exists. We recall that (ii) is equivalent to (i), and (i) is independent of the choice of $m$. In particular, if (ii) holds for a particular left-invariant mean, then (ii) holds for all left-invariant means, which shows that (ii) is equivalent to (v). Similarly, we see that (iii) is equivalent to (vi).
It is clear that (vii) implies (vi). To see that (vi) implies (vii), first note that, as above, the limit
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left\langle\pi_{s}\xi,x^{\prime}\right\rangle d\lambda(s)=d_{\mathcal{F}}-\displaystyle\lim_{s}\left\langle\pi_{s}\xi,x^{\prime}\right\rangle$
exists for any $\xi\in E$ and $x^{\prime}\in E^{\prime}$, so that $\underline{d}(\{s\in S:\pi_{s}\in U\})=1$ implies $d_{\mathcal{F}}(\{s\in S:\pi_{s}\in U\})=1$ for every weak neighbourhood of zero $U$ in $E$, and for any left-Følner net $\mathcal{F}$.
Now let us assume that $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence. It is now clear that (viii) implies (v), so we will proceed to show that (v) implies (viii). Since $(\pi_{s})_{s\in S}$ is uniformly bounded, we suppose without loss of generality (as mentioned in Section 2.4) that $S$ is a monoid and that $(\pi_{s})_{s\in S}$ consists of contractions. Then the adjoint action $(\pi_{s}^{*})_{s\in S}$ leaves the ($\sigma^{*}$-compact) dual unit ball $D^{\prime}\subseteq E^{\prime}$ invariant. Let $(T_{s})_{s\in S}$ denote the Koopman action on $C(D^{\prime})$ induced by $(\pi^{*}_{s})_{s\in S}$. For $\xi$ satisfying (ii) we let $f_{\xi}\in C(D^{\prime})$, $f_{\xi}(x^{\prime}):=\left|\langle\xi,x^{\prime}\rangle\right|^{p}$, $p>0$. We observe by assumption, that
@@
holds pointwise. Then, by the dominated convergence theorem the convergence of 14 holds weakly, i.e.,
@@
holds for any $\mu\in\left(C(D^{\prime})\right)^{\prime}$. By Eberlein’s generalized mean ergodic theorem, weak convergence implies the convergence of 15 in norm (see, e.g., Chapter 2, Theorem 1.5 of [32]). Hence
@@
Now let us show that (vi) is equivalent to (ix) under additional stated assumptions. Since $d_{\mathcal{F}}(C)=0$ for every compact set $C$, we see that (ix) implies (vi), so it remains to show that (vi) implies (ix). As in the previous case, we can suppose that every $\pi_{s}$ is a contraction. Since $E^{\prime}$ is separable, the weak topology on $E$ is first countable when restricted to the unit ball $E_{1}$, so let $\{U_{n}\}_{n=1}^{\infty}$ be a weak neighborhood basis for 0 in $E_{1}$. We can suppose it is decreasing, and to fix notations, set $U_{n}=V_{n}\cap E_{1}$ with $V_{n}$ weakly open in $E$. For each $n\in\mathbb{N}$, set $A_{n}=\{s\in S\ :\ \pi_{s}\xi\in U_{n}\}$; since every $\pi_{s}$ is contractive, we have also $A_{n}=\{s\in S\ :\ \pi_{s}\xi\in V_{n}\}$, so by (vi), $d_{\mathcal{F}}(A_{n})=1$. Let $m_{n}$ be such that $\lambda(A_{n}\cap F_{m})>\frac{n-1}{n}\lambda(F_{m})$ for all $m\geq m_{n}$. Assuming without loss of generality that $m_{n+1}>m_{n}$, we see that for $A:=\bigcup_{n=1}^{\infty}\bigcup_{m=m_{n}+1}^{m_{n+1}}(A_{n}\cap F_{m})$ we have $d_{\mathcal{F}}(A)=1$ and
$\{s\in A:\pi_{s}\xi\notin U_{n}\}=A\setminus A_{n}\subset\bigcup_{k<n}\;\bigcup_{m=m_{k}+1}^{m_{k+1}}(A_{k}\cap F_{m})\subseteq\bigcup_{m=1}^{m_{n}}F_{m}.$
@@
We now recall [30, Example 4.3] to show that Lemma 4.1 does not necessarily hold if $W(S)$ only admits a left-invariant mean and not a bi-invariant mean. This example will also show that Theorem 4.2 does not necessarily hold if $S$ is a left-amenable subsemigroup of an amenable group that is not right amenable.
We consider the semigroup $(S,\cdot)$ given by $S=\{(m,n)\in\mathbb{Z}^{2}\ :\ m\geq 0,n\geq 1\}$ with the multiplication
$(m_{1},n_{1})\cdot(m_{2},n_{2})=(2^{n_{2}}m_{1}+m_{2},n_{1}+n_{2}).$
As shown in [30], $S$ is cancellative and left-amenable, hence it satisfies the strong Følner condition, and since it is countable, it possesses a left-Følner sequence $\mathcal{F}=(F_{i})_{i=1}^{\infty}$. Furthermore, by replacing $F_{n}$ with $F_{n}\cdot(0,1)$ which form still a left Følner sequence, we may assume without loss of generality that $F_{n}\subseteq S\cdot(0,1)$ for all $n$. We observe that
$S\cdot(0,1)=\{(m,n)\in S\ :\ m\in 2\mathbb{Z},\;n\geq 2\},\quad S\cdot(1,1)=\{(m,n)\in S\ :\ m\notin 2\mathbb{Z},\;n\geq 2\}.$
Since $S\cdot(0,1)\cap S\cdot(1,1)=\emptyset$, we see that $S$ is not right amenable. Now let $\xi:S\rightarrow\{0,1\}$ be the characteristic function of $S\cdot(0,1)$. We note that for $t\in S$ and $s=(m,n)\in S$, we have $ts\in S\cdot(0,1)$ if and only if $m\in 2\mathbb{Z}$. We see that $\{R_{s}\xi\}_{s\in S}=\{0,1\}$, hence it is relatively weakly compact and $\xi\in W(S)$. Now set $E=W(S)$, $\pi=R$, and let $x^{\prime}$ denote the evaluation at $s_{0}$ for some $s_{0}\in S$. We see that $\xi\in E_{\text{aws}}$ because $R_{(1,1)}\xi=0$. However, for every $s\in F_{i}\subset S\cdot(0,1)$ we have $R_{s}\xi=1$, thus if $m\in\mathcal{M}(\mathcal{F})$, then
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|d\lambda(s)=m(s\mapsto\left(R_{s}\xi\right)(s_{0}))=m(1)=1.$ (16)
@@ 5 The JdLG decomposition via ultrafilters
Let $E=\mathcal{H}$ be a Hilbert space, $U$ a unitary operator on $\mathcal{H}$, and $\mathcal{H}=\mathcal{H}_{c}\oplus\mathcal{H}_{w}$ the compact-weak mixing decomposition. There are various results in the literature, such as [8, Theorem 2.2], [9, Theorem 2.25], [34, Theorem 4], and [11, Section 6], that (allow one to) construct the orthogonal projection $P:\mathcal{H}\rightarrow\mathcal{H}_{c}$ as an ultrafilter limit of powers of $U$. Theorem 5.2 is motivated by these previously mentioned results and attempts to generalize them to JdLG-admissible representations on a Banach space $E$. Strictly speaking, we cannot say that Theorem 5.2 generalizes all of the aforementioned results, because they also consider various classes of polynomial subsequences that we do not consider. For the next two results, we freely use the notation and facts from Subsection 2.5.
@@
We first prove (i). Let $h:\beta S\rightarrow\mathscr{S}_{r}=K(\mathscr{S})$ denote the map $h(q)=\displaystyle q-\lim_{s}\pi_{s}\xi$ with convergence taking place in the weak topology of $E_{r}$ (instead of the weak topology of $E$). Since $\pi:S\rightarrow\mathscr{S}$ is a homomorphism with dense image in a semitopological semigroup and $\pi=h\circ e$ on $S$, we see that $h$ a continuous surjective homomorphism. In particular, $h$ sends idempotents to idempotents. Since $\mathscr{S}_{r}=K(\mathscr{S})$ is a compact topological group, it contains a unique idempotent $Q$. Furthermore, we have already seen that $Q$ acts as the identity on $E_{r}$, so the desired result follows.
We now proceed to prove (ii). Theorem 4.2 tells us that for any weakly open neighborhood $U$ of 0 and every $\xi_{\text{aws}}\in E_{\text{aws}}$, we have
$d^{*}\left(\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\right)=0\text{, hence }\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\notin q.$ (17)
Since $q$ is an ultrafilter, we see that $\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\in U\right\}\in q$. Since $U$ was arbitrary, we see that $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$. ∎
@@
We begin with the proof of (i). We let $\mathscr{S}=c\ell_{\sigma}(\pi(S))$ and we recall that $\mathscr{S}$ is a compact semitopological semigroup. Let $\phi:\beta S\rightarrow\mathscr{S}$ be given by $\phi(q)=\pi_{q}=q-\lim_{s}\pi_{s}$, and observe that as in the proof of Lemma 5.1, $\phi$ is continuous surjective homomorphism. It follows that $K(\mathscr{S})=\phi(K(\beta S))$. Since $\pi$ is a JdLG-admissible representation, we see that $K(\mathscr{S})$ is a compact topological group with a unique idempotent $Q$, which is the projection onto $E_{r}$ along $E_{\text{aws}}$. Since $K(\mathscr{S})$ is closed, we see that $\phi\left(\overline{K(\beta S)}\right)=K(\mathscr{S})$. Since $\phi$ is a homomorphism, $\phi$ maps idempotents to idempotents, hence $\phi$ maps all idempotents of $\overline{K(\beta S)}$ to $Q$, which completes the proof of (i).